Find three consecutive integers whose sum is 33.
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52841)

.

The numbers are  (x-1), x  and (x+1).


The sum is  (x-1) + x + (x+1) = 33.


The equation is    3x = 33.


The solution is     x = 33/3 = 11.


The numbers are  10, 11, 12.    ANSWER


CHECK.  10 + 11 + 12 = 33.    ! Correct !

Solved.


Answer by Alan3354(69443)

Find three consecutive integers whose sum is 33.
-----------
33/3 is the middle integer.